# HySem: A context length optimized LLM pipeline for unstructured tabular extraction

### v1

@@ 2.3 Syntax Validator
Syntax errors in the LLM-generated JSON output render the table unusable for further table processing, such as analytics. Therefore, correcting syntax errors present in the JSON output is a critical functionality, particularly for facilitating automated workflows in industry settings. To address these challenges, we developed a Syntax Validator, based on reflective agentic framework.
Specifically, the Syntax Validator $\mathcal{A}_{\text{sv}}$ accepts a prompt $\mathcal{P}$ and a syntactically invalid JSON $\mathcal{J}_{\text{i}}$ as inputs and produces a syntactically valid JSON $\mathcal{J}_{\text{v}}$ through iterative refinement.
@@
Through self-reflection, $\mathcal{A}_{\text{sv}}$ iteratively refines the JSON output until a syntactically valid result is achieved or the maximum number of iterations is reached. The algorithm for this process is detailed in Algorithm 1.
Algorithm 1 Syntax Validator
@@
3 Evaluation Methodology
Building on this foundation, we develop intrinsic and extrinsic evaluation methods to measure the content and semantic accuracy of the JSON produced by HySem.
3.1 Intrinsic Evaluation
In intrinsic evaluation, we assess the representation of content from HTML cells in the generated JSON. We parse the HTML using Beautiful Soup to extract set of all the cell contents present in the HTML input, denoted as $\mathcal{H}_{\text{set}}=\{c_{1},c_{2},\ldots,c_{m}\}$. Similarly, we take the set of all elements in the JSON, denoted as $\mathcal{J}_{\text{set}}=\{e_{1},e_{2},\ldots,e_{n}\}$
The Intrinsic Score (ISC) is then computed as:
$\text{ISC}=\frac{\sum_{i=1}^{m}I(c_{i}\in\mathcal{J}_{\text{set}})}{m}.$
Here, $m$ denotes the cardinality of $\mathcal{H}_{\text{set}}$.
3.2 Extrinsic Evaluation
$\mathcal{P}=\{p_{i}\mid p_{i}=(n_{1},n_{2},\ldots,n_{k})\}$
Here, $(n_{1},n_{2},\ldots,n_{k})$ represents a sequence of nodes starting from the root node $n_{1}$ to a leaf node $n_{k}$ in the JSON structure. For each path $p_{i}\in\mathcal{P}$, we prompt an LLM ($\mathcal{M}_{\text{q}}$) with $\mathcal{G}_{\text{t}}$ and $p_{i}$ as inputs, instructing it to generate a single question $\mathcal{Q}_{\text{i}}$. This question is targeted so that the value at the leaf node of the path is the expected answer $\mathcal{K}_{\text{e}}$.
The question-answer pairs are evaluated using an Evaluator LLM ($\mathcal{M}_{\text{eval}}$). This model accepts HySem generated JSON $\mathcal{J}_{\text{p}}$, question $\mathcal{Q}_{\text{i}}$ and an expected answer $\mathcal{K}_{\text{e}}$ as inputs. $\mathcal{M}_{\text{eval}}$ predicts an answer $\mathcal{K}_{\text{p}}$ and compares its prediction with the expected answer $\mathcal{K}_{\text{e}}$ in a single pass.
$Pred(p_{i})=\left\{\begin{array}[]{ll}1&\text{if }\mathcal{K}_{\text{p}}=\mathcal{K}_{\text{e}},\\ 0&\text{otherwise.}\end{array}\right.$
The Extrinsic Score (ESC) is computed as:
$ESC=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}\text{Pred}(p_{i}).$
4 Results and Ablation Study
We present the results of HySem following the first fine-tuning stage, utilizing our 608 testing samples manually curated from the open-sourced FinTabNet and PubTabNet datasets. This comparison highlights HySem’s performance after being fine-tuned on publicly available data, ensuring a fair and transparent benchmark against other models.
Model Intrinsic Score % Extrinsic Score %
Meta-Llama-3-8B-Instruct 66 84.42
Phi-3-Medium-128K-Instruct 56 82.39
GPT-4o 93.4 90.15
HySem 91.12 88.39
Table 3: Intrinsic and Extrinsic Evaluation results
We benchmarked our model against popular open-source models, Meta-Llama-3-8B-Instruct and Microsoft’s Phi-3-Medium-128K-Instruct, as well as GPT-4o. When compared against Llama 3, HySem acheives over 25% higher accuracy on the intrinsic score and 3.97% higher on the extrinsic score. On the same test dataset, Microsoft Phi-3 Medium model performed poorly on the intrinsic metrics and fared relatively better at 82.39% on the extrinsic benchmark.
Number of test samples 608
Number of tokens before optimization (A) 570330
Number of tokens after optimization (B) 233287
Average number of tokens before optimization 938
Average number of tokens after optimization 383
Token Efficiency(%) 59.09
Table 4: Token efficiency metrics
We evaluated the token reduction efficiency of the Context Optimizer. Our studies as indicated by Table 4., show impressive reductions in the token count resulting in a reduced LLM inference throughput. We define Token Efficiency as
The results show that HySem significantly outperforms both the open-source models that we evaluated and remains competitive with the industry-leading OpenAI GPT-4o in both intrinsic and extrinsic measures with a token efficiency of over 59%. Although our model trails GPT-4o by 2% in accuracy metrics, it offers substantial advantages such as on-premise deployment on commodity hardware, cost-effectiveness, and enhanced data privacy—non-negotiable requirements for many enterprises.
5 Related Works
Recent research has explored various uses of HTML to enhance NLP tasks. Aghajanyan et al. (2021) developed the HyperText Language Model (HTLM), which leverages HTML’s structural elements to improve tasks like zero-shot summarization and question answering. Their approach utilizes HTML for structured prompting and template-based guidance but is largely confined to standard NLP tasks and assumes predefined HTML structures, such as <title> elements.
Gur et al. (2023) focused on HTML understanding through tasks like semantic classification, description generation, and autonomous web navigation. They identified the context window length as a significant bottleneck, noting that even models with support for 1000+ tokens struggle with performance degradation when processing larger snippet sizes. This challenge underscores the difficulty of managing extensive HTML content within fixed context constraints.
Sui et al. (2024) tackled context length limitations by serializing tables into text formats. Their strategies include varying input formats and identifying critical values through self-augmentation. While these methods help fit table data within the model’s context window, serialization can lead to the loss of structural nuances and detailed information.
In contrast, our work is broadly categorized as a Table Transformation task, where we focus on converting arbitrary HTML tables into semantic JSON. By employing a novel Context Optimizer, we preserve the table’s structural integrity while enabling the processing of large tables. Unlike methods relying on predefined templates or serialization, our approach does not make assumptions about the HTML structure.
6 Conclusion
We successfully implemented the HySem LLM Pipeline, which generates semantic JSON output from HTML tables. The Context Optimizer significantly improved context utilization, enabling us to process larger tables. Currently, we are piloting our product with a well-known pharmaceutical enterprise that operates multiple production plants and serves a global customer base.
This model powers several downstream applications, including analytics from regulatory compliance documents and the automatic creation of these documents.
Our future work includes supporting even larger tables that span multiple pages, developing techniques to improve processing speed, and building custom models for other verticals.

### v2

@@ 2.3 Syntax Corrector
Syntax errors in the LLM-generated JSON output render the table unusable for further processing, such as ingestion into databases. Consequently, correcting these syntax errors is a critical functionality, especially for enabling automated workflows in industrial settings. To address these challenges, we developed a Syntax Corrector, based on reflective agentic framework.
@@
Specifically, the Syntax Corrector $\mathcal{A}_{\text{sc}}$ accepts a syntactically invalid JSON $\mathcal{J}_{\text{i}}$ as input and produces a syntactically valid JSON $\mathcal{J}_{\text{v}}$ through iterative refinement. Through self-reflection (Ji et al., 2023; Pan et al., 2024a; Madaan et al., 2024; Asai et al., 2023; Renze and Guven, 2024), $\mathcal{A}_{\text{sc}}$ iteratively refines the JSON output until a syntactically valid result is achieved or the maximum number of iterations is reached. The algorithm for this process is detailed in Algorithm 2 of Appendix A.5. Table 2 displays the common syntax error patterns observed in the LLM output.
3 Evaluation Methodology
Building on this foundation, we develop intrinsic and extrinsic evaluation methods to measure the content and semantic accuracy of the JSON produced by HySem. Appendix A.4 provides detailed illustrations for intrinsic and extrinsic evaluation methods.
3.1 Intrinsic Evaluation
In intrinsic evaluation, we assess the representation of content from HTML cells in the generated JSON. We parse the HTML using Beautiful Soup to extract the set of all the cell contents present in the HTML input, denoted as $\mathcal{H}_{\text{set}}=\{c_{1},c_{2},\ldots,c_{m}\}$. Similarly, we take the set of all elements in the JSON, denoted as $\mathcal{J}_{\text{set}}=\{e_{1},e_{2},\ldots,e_{n}\}$
The Intrinsic Score (ISC) is then computed as:
$\text{ISC}=\frac{\sum_{i=1}^{m}I(c_{i}\in\mathcal{J}_{\text{set}})}{m}.$
Here, $m$ denotes the cardinality of $\mathcal{H}_{\text{set}}$.
3.2 Extrinsic Evaluation
$\mathcal{P}=\{p_{i}\mid p_{i}=(n_{1},n_{2},\ldots,n_{k-1})\}$
Here, $(n_{1},n_{2},\ldots,n_{k})$ represents a sequence of nodes starting from the root node $n_{1}$ till a leaf node $n_{k}$ in the JSON structure. For each path $p_{i}\in\mathcal{P}$, we prompt an LLM ($\mathcal{M}_{\text{q}}$) with $\mathcal{G}_{\text{t}}$ and $p_{i}$ as inputs, instructing it to generate a single question $\mathcal{Q}_{\text{i}}$. This question is targeted so that the value at the leaf node of the path is the expected answer $\mathcal{K}_{\text{e}}$.
An Evaluator LLM ($\mathcal{M}_{\text{eval}}$) takes as input the HySem-generated JSON ($\mathcal{J}_{\text{p}}$), a question ($\mathcal{Q}_{\text{i}}$), and the expected answer ($\mathcal{K}_{\text{e}}$). The Evaluator LLM predicts an answer ($\mathcal{K}_{\text{p}}$) and compares its prediction with the expected answer ($\mathcal{K}_{\text{e}}$), all in a single pass through the LLM.
Hypothesis: Verifying each leaf node in the JSON is sufficient to confirm the structure’s accuracy. If the JSON is incorrect, it will fail to answer questions about these leaf nodes accurately.
The Evaluator LLM computes the score for the $i$-th question-answer pair as follows:
$\mathcal{M}_{\text{eval}}(p_{i})=\left\{\begin{array}[]{ll}1&\text{if }\mathcal{K}_{\text{p}}=\mathcal{K}_{\text{e}},\\ 0&\text{otherwise.}\end{array}\right.$
The Extrinsic Score (ESC) is computed as:
$ESC=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}\text{Pred}(p_{i}).$
4 Results
We present the results of HySem, measured on 608 testing samples manually annotated from the open-sourced FinTabNet and PubTabNet datasets. The performance of our pipeline was evaluated using both intrinsic and extrinsic metrics, as described in the previous section, ensuring a thorough and well-rounded assessment. The results demonstrate that HySem delivers competitive performance compared to industry-leading models and outperforms popular open-source LLMs while excelling significantly in terms of token efficiency.
4.1 Baselines and Benchmarking
Our method uses LLaMA-3-Base as the foundational LLM, which we fine-tune to enhance performance for our specific table transformation task. To assess the improvements gained, we benchmarked HySem against popular open-source and proprietary models, including Meta LLaMA-3-8B-Instruct, Microsoft Phi-3-Medium-128K-Instruct, and OpenAI GPT-4o. Table 3 shows our benchmarking results.
LLaMA-3-8B-Instruct: HySem surpasses LLaMA-3-8B-Instruct by over 25% in intrinsic accuracy and 3.97% in extrinsic scores, proving that our task-specific fine-tuning significantly enhances the model’s capabilities.
Phi-3-Medium-128K-Instruct: The Phi-3 model displayed weak performance on intrinsic metrics but fared better in extrinsic evaluations, achieving 82.39%. HySem outperformed Phi-3 in both areas, demonstrating its superiority as a fine-tuned model specialized for this task.
GPT-4o: While GPT-4o demonstrates superior overall accuracy metrics, largely due to extensive training on diverse datasets, particularly involving HTML and XML table formats (Sui et al., 2024), HySem maintains a competitive edge. Although trailing by 2% in accuracy, HySem offers substantial advantages such as on-premise deployment on commodity hardware, cost-effectiveness, and enhanced data privacy—essential requirements for many enterprises.
Model Intrinsic Score Extrinsic Score
Meta-Llama-3-8B-Instruct 66.25 84.42
Phi-3-Medium-128K-Instruct 56.78 82.39
GPT-4o 93.43 90.15
HySem (Ours) 91.12 88.39
Table 3: Benchmarking Results
Description Metric
Number of test samples 608
Number of tokens before Token-based Encoding (A) 366608
Number of tokens after Token-based Encoding (B) 224082
Token Efficiency(%) 38.87
Table 4: Token Efficiency Metrics
4.2 Token Reduction Efficiency
We evaluated the token reduction efficiency of HySem’s Context Optimizer, which enhances token usage while maintaining semantic accuracy. HySem improves model efficiency, resulting in faster inference times and reduced computational overhead.
Token Efficiency is defined as:
Table 4 shows that HySem achieves over 38% token efficiency, reflecting a substantial reduction in the token count required for processing, thereby enhancing the LLM inference throughput. The results demonstrate significant benefits gained with using our Token-based encoding method.
5 Conclusion and Future Work
We successfully implemented the HySem LLM Pipeline, which generates semantic JSON output from HTML tables. The Context Optimizer significantly improved context utilization, enabling us to process larger tables. Currently, we are piloting our product with a well-known pharmaceutical enterprise that operates multiple production plants and serves a global customer base. This model powers several downstream applications, including analytics from regulatory compliance documents and the automatic creation of these documents. Our future work includes supporting even larger tables that span multiple pages, developing techniques to improve processing speed, and building custom models for other verticals.
Appendix A Appendix
A.1 Related Works
LLMs for Tabular Data
Tables are a fundamental method for presenting structured information across various industrial applications, including regulatory information in the pharmaceutical industry and financial reports for businesses. Some shared characteristics and inherent challenges in tables as mentioned by [Fang et al., 2024] include Heterogeneity [Borisov et al., 2023], Sparsity, Context-based interconnection [Liu et al., 2023], Lack of prior knowledge [Borisov et al., 2022, 2023], etc. Recent advancements in Large Language Models [Achiam et al., 2023, Jiang et al., 2023, Team et al., 2024, Dubey et al., 2024] have shown promise in addressing these challenges for various table related tasks. Some table related tasks and techniques include entity matching and data-imputation [Li et al., 2024, Korini and Bizer, 2023], tabular Q&A [Chen et al., 2024, Wang et al., 2024, Zhang et al., 2023, Patnaik et al., 2024], schema augmentation [Narayan et al., 2022], Serialization [Yin et al., 2020, Yan et al., 2024], Table Manipulation [Zhang et al., 2024], Table Understanding [Lobo et al., 2023], Prompt Engineering [Zheng et al., 2023, Ziqi and Lu, 2023], Table RAG [Sundar and Heck, 2023]. Role-play [Zhao et al., 2023].
LLMs and HTML
Recent research has explored various uses of HTML to enhance NLP tasks. [Nakano et al., 2021, Yao et al., 2022] utilize LLMs for Web Navigation tasks involving HTML input formats. [Aghajanyan et al., 2022] developed the HyperText Language Model (HTLM), which leverages HTML’s structural elements to improve tasks like zero-shot summarization and question answering. Their approach utilizes HTML for structured prompting and template-based guidance but is largely confined to standard NLP tasks and assumes predefined HTML structures, such as <title> elements. [Gur et al., 2023] focused on HTML understanding through tasks like semantic classification, description generation, and autonomous web navigation. They identified the context window length as a significant bottleneck, noting that even models supporting longer token sequences struggle with performance degradation when processing larger snippet sizes.
In contrast to the aforementioned works, our approach focuses not merely on HTML, but specifically on the more complex domain of HTML tables particularly Pharmaceutical tables, which present significant challenges due to their hierarchical and relational structure. HTML tables often contain multi-level headers, rowspan and colspan attributes, and intricate relationships between cell elements, making their transformation into a semantic format far more demanding. Our work uniquely tackles this complexity by performing a table transformation task, converting HTML tables into semantic JSON. This transformation is further enhanced by our Context Optimizer, which efficiently reduces token usage while preserving the intricate relationships between table elements. Unlike previous approaches, which focus on predefined HTML structures, our method handles the complexities of tables dynamically, ensuring both structural accuracy and computational efficiency. To the best of our knowledge, this is the first approach to transform complex HTML tables into semantic JSON in a context-length optimized manner.
Context Optimization for Tables in LLMs
Recent advancements in context-length optimization have primarily focused on improving LLM performance for text-based tasks [Pan et al., 2024b, Fountas et al., 2024, Han et al., 2024], but limited work has been done to address the unique challenges presented by tables. [Sui et al., 2024] tackled context length limitations by serializing tables into text formats. While these methods help fit table data within the model’s context window, serialization can lead to the loss of structural nuances and detailed information. In contrast, [Tian et al., 2024] introduces a novel encoding method tailored for spreadsheet data. While effective in reducing token usage and computational costs, it addresses a different use-case (spreadsheets) compared to our work (HTML tables). Additionally, the encoding method primarily leverages format and structure based optimisations, whereas our approach rewrites the input to ensure alignment with the tokenizer. Our approach specifically targets HTML tables of arbitrary structure, ensuring that hierarchical and relational elements are preserved during HTML to JSON transformation task. By utilizing our Context Optimizer, we ensure semantic integrity is maintained even during token pruning, unlike text-based methods where reducing detail (e.g., abbreviating "potassium clavulanate" to "potassium") would alter the meaning. This is a significant advantage in transforming HTML tables into semantic JSON without compromising accuracy.
A.2 Complex Pharmaceutical Data Table
Figure 3: This table compares various dosage regimens of Glucophage and Glucophage XR. The challenge in converting it to semantic JSON lies in its nested categories, including dosage, measurement types, and time points. Each combination of dose, metric, time point, and statistical details (mean change, confidence intervals) must be accurately mapped to database schema keys. Ensuring data integrity while managing variability in column structures is essential for creating a usable semantic representation.
A.3 Semantic JSON
In our paper, the term "Semantic JSON" refers to a JSON representation where the keys are designed to accurately mirror the hierarchical structure of the JSON tree and align directly with the database schema. This ensures that the JSON output is not only well-structured but also semantically meaningful. Figure 7 provides an example of such a Semantic JSON output, as generated by Hysem. To highlight the effectiveness of our approach, we compare the outputs generated by other models, namely Llama 3 8B Instruct and Microsoft Phi 3 128K Instruct, as illustrated in Figures 7 & 7 respectively. These illustrations show how the JSON representations from these models differ from ours, particularly in terms of hierarchical accuracy and schema alignment.
Figure 4: Sample Table to illustrate Semantic JSON
HySem output versus Meta LLama 3 8B Instruct and Microsoft Phi 3 128K Instruct
The sample illustrates that the Llama 3 output has missed certain fields, such as ’Slipped Side’ and ’Non-slipped Side.’ In contrast, Phi 3 has introduced several errors, including misplacing values under the incorrect hierarchy. For instance, the p-value of 0.717 is erroneously placed under the ’range_values’ subtree for the ’Non-slipped Side.’
Figure 5: Hysem JSON
Figure 6: LlaMa3 JSON
Figure 7: Phi3 JSON
A.4 Evaluation Methodology Illustrations
The performance of HySem is assessed using both intrinsic and extrinsic evaluation methods. Intrinsic evaluation focuses on how accurately the input HTML strings are converted to JSON, while extrinsic evaluation assesses how effectively the semantic structure is preserved.
Figure 8: Sample of HTML Table for illustration
For the given HTML table, the HySem generated semantic JSON and the ground-truth (GT) JSON are as below.
Figure 9: Hysem JSON
Figure 10: GT JSON
The Intrinsic Evaluator’s evaluation results are presented in Table 5. Here, each unique cell content of the original HTML table is compared directly to the nodes of the JSON.
We initially used metrics to measure how well the JSON preserved the "frequency" of each string from the HTML. However, this method presented problems. For e.g., a cell content could appear repeatedly in the JSON for each row entry, unlike its single occurrence in the HTML. This discrepancy can skew the accuracy of metrics based on string frequency.
To address this, we simplified the approach by ensuring each unique string from the HTML appears at least once in the JSON. While it’s a reasonable estimate of performance, we recognise that this is an area for future improvement.
Cell Content Presence in Hysem JSON (Yes/No)
"Gum use"
"Time"
"Baseline"
"Polyol"
"Subjects (n)"
"90"
"Mean ± SD"
"5.32 ± 0.43"
"Xylitol"
"89"
"5.41 ± 0.35"
"p value one-way ANOVA"
"0.29"
"6 months"
"79"
"5.22 ± 0.21"
"77"
"5.33 ± 0.46"
"0.31"
"12 months"
"72"
"5.33 ± 0.42"
"71"
"5.16 ± 0.42"
"0.03"
"No-gum use"
"24 months"
"64"
"5.33 ± 0.46"
"66"
"5.15 ± 0.64"
"0.04"
"<0.01" x
"0.42" x
Intrinsic Score: 94.11%
Table 5: Intrinsic Evaluation results
The Extrinsic evaluation measures the semantic accuracy of the LLM generated JSON. In the example table, there are 21 unique paths from the root node to each leaf node of the GT JSON. As mentioned, the number of paths is equal to the number of leaf nodes. LLM ($\mathcal{M}_{\text{q}}$) accepts GT JSON and a path and outputs a single question such that the expected answer is the leaf node of the path. Table 6 illustrates the paths, corresponding LLM generated questions and the GT answer.
Path Question GT Answer
Gum use > Time > Baseline > Polyol > Subjects (n) What is the number of subjects in the baseline group that used gum containing polyol? 90
Gum use > Time > Baseline > Polyol > Mean ± SD What is the average amount of gum used by subjects at baseline, with a standard deviation? 5.32 ± 0.43
Gum use > Time > Baseline > Xylitol > Subjects (n) What is the number of subjects in the baseline group that used Xylitol? 89
Gum use > Time > Baseline > Xylitol > Mean ± SD What is the average amount of Xylitol used by subjects at baseline, with a standard deviation? 5.41 ± 0.35
Gum use > Time > Baseline > p value one-way ANOVA What is the p-value of the one-way ANOVA test for the baseline data? 0.29
Gum use > 6 months > Polyol > Subjects (n) What is the number of subjects in the group that used gum for 6 months and was given polyol? 90
Gum use > 6 months > Polyol > Mean ± SD What is the mean value of gum usage after 6 months for polyol? 5.22 ± 0.21
Gum use > 6 months > Xylitol > Subjects (n) What is the number of subjects in the group that used Xylitol for 6 months? 77
Gum use > 6 months > Xylitol > Mean ± SD What is the mean value of Xylitol at 6 months, along with its standard deviation? 5.33 ± 0.46
Gum use > 6 months > p value one-way ANOVA What is the p-value of the one-way ANOVA test for the comparison between Polyol and Xylitol at 6 months? 0.31
Gum use > 12 months > Polyol > Subjects (n) How many subjects were in the group that used gum for 12 months and had a polyol treatment? 72
Gum use > 12 months > Polyol > Mean ± SD What is the average value of gum use for 12 months with polyol, along with its standard deviation? 5.33 ± 0.42
Gum use > 12 months > Xylitol > Subjects (n) What is the number of subjects in the group that used Xylitol for 12 months? 71
Gum use > 12 months > Xylitol > Mean ± SD What is the mean value of Xylitol at 12 months, along with its standard deviation? 5.16 ± 0.42
Gum use > 12 months > p value one-way ANOVA What is the p-value of the one-way ANOVA test for the comparison at 12 months? 0.03
No-gum use > 24 months > Polyol > Subjects (n) How many subjects were in the group that did not use gum and had a follow-up at 24 months, with a focus on polyol? 64
No-gum use > 24 months > Polyol > Mean ± SD What is the mean value of gum usage after 24 months for subjects using polyol, along with its standard deviation? 5.33 ± 0.46
No-gum use > 24 months > Xylitol > Subjects (n) How many subjects were in the group that did not use gum and had data collected at 24 months, with a focus on xylitol? 66
No-gum use > 24 months > Xylitol > Mean ± SD What is the mean value of Xylitol at 24 months for subjects with no gum use, along with its standard deviation? 5.15 ± 0.64
No-gum use > 24 months > p value one-way ANOVA > Polyol > Mean ± SD What is the mean difference in standard deviation of ’p value one-way ANOVA’ for ’Xylitol’ for ’Polyol’ at 24 months in a group with no gum use? 0.42
No-gum use > 24 months > p value one-way ANOVA > Xylitol > Mean ± SD What is the mean difference in standard deviation of ’p value one-way ANOVA’ for ’Xylitol’ at ’24 months’ under ’No-gum use’? < 0.01
Table 6: Paths, Questions and GT Answers of Extrinsic Evaluation
The evaluator LLM ($\mathcal{M}_{\text{eval}}$), predicts an answer, given the LLM-generated JSON, a single question, and the expected answer as inputs. It then compares its prediction with the expected answer to generate a score, all in a single pass through the LLM. Table 7 shows the predicted answers and the extrinsic scores computed for each question, along with the overall extrinsic score.
Predicted Answer Score
90 1
5.32 ± 0.43 1
89 1
5.41 ± 0.35 1
0.29 1
90 1
5.22 ± 0.21 1
77 1
5.33 ± 0.46 1
0.31 1
72 1
5.33 ± 0.42 1
71 1
5.16 ± 0.42 1
0.03 1
64 1
5.33 ± 0.46 1
66 1
5.15 ± 0.64 1
0.04 0
0.03 0
Extrinsic Score 90.47%
Table 7: Extrinsic Evaluation results
A.5 Syntax Corrector: Algorithm
Our Semantic Synthesizer incorporates a custom fine-tuned model specialized in generating JSON outputs. However, occasional syntax errors may occur in the generated output. The Syntax Corrector, part of the HySem pipeline, handles these erroneous JSONs by performing LLM-assisted auto-correction via self-reflection. We instruct the LLM with the following prompt:
"The JSON given below contains syntax errors. Your task is to correct them and provide the corrected output. Provide ONLY the corrected output, without any additional explanation. ’input_string’: {json_input}"
Algorithm 2 presents the syntax correction algorithm.
Algorithm 2 Syntax Corrector